I) a (f) b = a/b – 2(b f a)
II) 3 (f) 1 = -7/9
III) 2 (f) 1 = ?
In the equations above, the function of (f) remains unchanged. What does the question mark stand for in the third equation?
a) -2/3, b)-1/3, c) 2/3, d)1, e)3

Found 2 solutions by ikleyn, josgarithmetic:
Answer by ikleyn(52778)

.
You use absolutely non-standard symbols/denotations for a (the) function without any explanations, which are necessary and are supposed to be.

Do you expect I will read between lines ??


Find another site.


Or write in a way accepted in Math.



Answer by josgarithmetic(39617)

The purpose for (I) must have been to show how the OPERATION (f) is supposed to work.

The purpose of (II) seems unnecessary other than to help to be sure we have used the operation as defined in (I) correctly.

Following the definition for f demonstrated in (I), choice (b) seems to be the result for (2 f 1); .

III.
2f1=2/1-2(1f2)
=2-2(1/2-2(2f1))
=2-1+4(2f1)
2f1=2-1+4(2f1)
2f1=1+4(2f1)
0=1+3(2f1)
-1=3(2f1)
-1/3=2f1
Corresponding to choice b.